# Debiased Nonparametric Regression for Statistical Inference and Distributionally Robustness

### v2

@@ 5 Conclusion
This study developed a debiased estimator for nonparametric regression. For any smooth regression estimator, our debiased estimator gives the following desirable property: mean-squared convergence, asymptotic normality, uniform convergence, and double robustness. The asymptotic normality is useful for statistical inference, while the uniform convergence makes the nonparametric regression estimator distributionally robust. Additionally, either the first-stage or the second-stage nonparametric estimator is consistent, the resulting regression estimator is consistent; that is, doubly robust. Surprisingly, our estimator only requires the smoothness of the first-stage regression estimator. More specifically, if either (i) both the first-stage estimator $\widehat{f}$ and the true regression function $f_{0}$ are smooth or (ii) the difference $\widehat{f}-f_{0}$ is smooth, we can show the above desirable properties for the debiased estimator.
6 Proofs
@@
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$ $\coloneqq{\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}.$
@@
$\mathrm{Variance}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right]+\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right).$
We aim to show that
@@
$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)=\frac{C_{2}}{nh^{d}_{n}}+o(1),$ (3)
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)=2C^{2}_{1}h^{2s}.$ (4)
@@
$=\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u,$
@@
$\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u$
$=L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\int\left|\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(u\right)\mathbbm{1}\left[\|u\|\leq 1\right]\right|{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u.$
6.1.1 Preliminaries
@@
Lemma 6.1 (From Proposition 1.12 in Tsybakov (2008)).
@@
Lemma 6.2.
Under Assumption 3.2, for any $\varepsilon$, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h\geq\frac{1}{2n}$, and $x_{0}\in\mathcal{X}$, the weights $\widehat{w}_{n}(x,x_{0})$ of the $\ell$) estimator satisfy the followings:
@@
6.1.2 Bounding the bias (proof of (2))
@@
Let $\widehat{g}_{n}\coloneqq\widehat{f}_{n}-f_{0}$. Since $\widehat{g}_{n}$ belongs to the Hölder class $\Sigma(s,L)$, we have
@@
where $0\leq\tau_{i}\leq 1$. Here, we used the Taylor expansion of $\widehat{g}$, and Lemma 6.1.
With probability one, we have
@@
6.1.3 Bounding the variance: part I (proof of (3))
@@
We have
@@
From Assumption 3.2, we have
@@
$\leq\frac{\sigma^{2}_{\max}C^{2}_{*}}{mh^{d}_{n}}+o(1)$
$=\frac{C_{2}}{nh^{d}_{n}}+o(1).$
6.1.4 Bounding the variance: part II (proof of (4))
@@
We have
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
@@
We decompose the sup-norm as follows:
@@
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh^{d}_{n}},$
@@
We have
@@
$h_{n}=\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}},$

### v3

@@ 6 Conclusion
This study developed a debiased estimator for nonparametric regression. For any smooth regression estimator, our debiased approach ensures pointwise and uniform MSE convergence, asymptotic normality, and double robustness. The asymptotic normality facilitates statistical inference, while uniform convergence provides distributional robustness for nonparametric regression. Additionally, if either the first-stage or the second-stage nonparametric estimator is consistent, then the resulting regression estimator remains consistent, exhibiting the doubly robust property.
Notably, our estimator requires only the smoothness of the first-stage regression estimator. Specifically, if either (a) both the first-stage estimator $\widehat{f}$ and the true regression function $f_{0}$ are smooth, or (b) the difference $\widehat{f}-f_{0}$ is smooth, then the desirable properties of the debiased estimator hold. Thus, our method imposes minimal structural constraints while guaranteeing strong theoretical results, making it an attractive and flexible approach for a wide range of nonparametric regression problems.
Appendix
Appendix A Proofs
@@
$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \coloneqq{\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}.$
@@
$\mathrm{Variance}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right]+\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right).$
We aim to show that the followings hold with probability one:
@@
$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)=\frac{C_{2}}{nh_{n}}+o(1),$ (3)
$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)=2C^{2}_{1}h^{2s}.$ (4)
@@
$=\int L\frac{h^{s}_{n}}{\ell!}h_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u,$
@@
$\int L\frac{h^{s}_{n}}{\ell!}h_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u$
$=L\frac{h^{s}_{n}}{\ell!}h_{n}\int\left|\frac{1}{h_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(u\right)\mathbbm{1}\left[\|u\|\leq 1\right]\right|{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u.$
A.1.1 Preliminaries
@@
Lemma A.1 (From Proposition 1.12 in Tsybakov (2008)).
@@
Lemma A.2.
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
  • There exists a constant $C>0$ such that
    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
    where $P(x)$ is the distribution of $x$.
For any $\varepsilon$, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h\geq\frac{1}{2n}$, and $x_{0}\in\mathcal{X}$, the weights $\widehat{w}_{n}(x,x_{0})$ of the $\ell$) estimator satisfy the followings:
@@
A.1.2 Bounding the bias (proof of (2))
@@
Let $\widehat{g}_{n}\coloneqq\widehat{f}_{n}-f_{0}$. Since $\widehat{g}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, we have
@@
as $n\to\infty$, where $0\leq\tau_{i}\leq 1$. Here, we used the Taylor expansion of $\widehat{g}$, and Lemma A.1.
Therefore, the followng holds almost surely as $n\to\infty$:
@@
A.1.3 Bounding the variance: part I (proof of (3))
@@
As $n\to\infty$, the following holds almost surely:
@@
From the conditions we assumed, as $n\to\infty$, the following holds:
@@
$\leq\frac{\sigma^{2}_{\max}C^{2}_{*}}{mh_{n}}+o(1)$
$=\frac{C_{2}}{nh_{n}}+o(1).$
A.1.4 Bounding the variance: part II (proof of (4))
@@
As $n\to\infty$, the following holds t
$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
@@
As $n\to\infty$, the following holds almost surely:
@@
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh_{n}},$
@@
As $n\to\infty$, the following holds almost surely:
@@
$h_{n}=\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+1}},$